The following information is about a Company

Month Sales Purchases

Jan. 60,000 30,000

Feb. 80,000 40,000

Mar. 100,000 60,000

Cash is collected from customers in the following manner:

Month of sale 45%

Month following the sale 60%

40% of purchases are paid for in cash in the month of purchase, and the balance is paid in the following month.

Labor costs are 30% of sales. Other operating costs are 38,000 per month including deprecation of 10,000. Both of these are paid in the month incurred.

How much cash will be disbursed in total in March?
Answer #1
Payment for February purchases 24000 =40000*60%
Payment for March purchases 24000 =60000*40%
Payment for labor costs 30000 =100000*30%
Payment for operating costs 28000 =38000-10000
Cash disbursed in total in March 106000